OpenStudy (anonymous):

Find the number of real solutions using the discriminant. x² + 9x + 4 = 0

OpenStudy (anonymous):

Calculate b^2-4ac

OpenStudy (anonymous):

2

OpenStudy (campbell_st):

this is an interesting question as all quadratics will have 2 solutions. b^2 - 4ac > 0 then the quadratic has 2 real unequal solutions if the value of b^2 - 4ac is a square number then the solutions are real and rational. if b^2 - 4ac is not a square number the solution is irrational. if b^2 - 4ac = 0 then the quadratic has a repeated or 2 equal solution. This is the case of the perfect square y = (x -a)^2 in your question you have 9^2 - 4 x 1 x 4 = 81 - 16 the discriminant = 65 there are 2 real irrational solutions